# Algebraic Robustness Verification of Neural Networks

### v1

@@ 3.1 ED Degree for Generic Parameters
Our framework allows us to distinguish between the geometric complexity of the boundary and the identifiability of the parameters. Let $p$ be the number of parameters and $N$ be the dimension of the space of degree-$D$ polynomials. Let $\Phi:\mathbb{R}^{p}\to\mathbb{R}^{N}$ be the parameterization map sending weights to coefficients. The observed drop in ED degree is governed by the dimension of the image $\mathrm{Im}(\Phi)$ relative to the ambient dimension $N$. For intermediate architectures, we typically have $\dim(\mathrm{Im}(\Phi))\ll N$. This strict inequality ensures that the decision boundary lies on a proper subvariety of the polynomial space, which explains the reduced algebraic complexity. In contrast, parameter identifiability is determined by the dimension of the image relative to the domain dimension $p$. If $\dim(\mathrm{Im}(\Phi))<p$, then by the Fiber Dimension Theorem (Hartshorne, 1977, Chapter 1), the fibers $\Phi^{-1}(f)$ are positive-dimensional varieties, and thus the network is defective. This comparison clarifies the trade-off: the ED degree depends on how small the image is inside $\mathbb{R}^{N}$, while identifiability depends on how compressed the map is from $\mathbb{R}^{p}$. A network can thus achieve low verification complexity while remaining non-identifiable if the architecture is over-parameterized.
@@ 3.2 Expected Real ED Degree
The results in the preceding section focus on critical points of the distance minimization problem over the complex field. For computation with homotopy continuation this is indeed the relevant quantity to consider. However, in some cases one may want to focus on the real critical points only. In practice, the number of real ED critical points is typically much smaller than the ED degree and may vary across different generic parameter realizations. To capture this behavior, we study the expected real ED degree, for a distribution of parameters. We leverage the Kac–Rice formalism (Kac, 1943; Rice, 1944) to derive a formula for this expectation, quantifying the typical number of real critical points encountered in robustness verification.
@@
This gives the expected number of real critical points of the distance function from $\bm{\xi}$ to the random hypersurface defined by $f(\bm{x})=0$ within the set $U$. The formula is very general and necessarily involved. We can abstract away some of the details by considering NNGPs (Lee et al., 2018):
Corollary 3.11.
Let $f:\mathbb{R}\to\mathbb{R}$ be the neural network Gaussian process (NNGP) induced by an infinite-width two-layer network with polynomial activation of degree $d\geq 1$. Let $U=\mathbb{R}\times(\mathbb{R}\setminus\{0\})$. Then the expected real ED degree is
$\mathbb{E}\Big[\#\{\,(x,\lambda)\in U:F(x,\lambda)=0\,\}\Big]=\frac{d}{\sqrt{2d-1}}.$
Remark 3.12.
In this one-dimensional case, by Proposition˜3.2, an infinite-width two-layer network with degree-$d$ activation has ED degree $d$, whereas on average only $\frac{d}{\sqrt{2d-1}}$ of these critical points are real. In contrast, for Kac polynomials of degree $d$, the expected number of real roots is about $O(\log d)$ (Breiding and Lerario, 2024). Thus, the special algebraic structure induced by the network parametrization significantly increases the expected number of real critical points compared to Kac polynomials.
@@ 4.1 ED Discriminant
The notion of ED discriminant was introduced in the context of algebraic optimization by Draisma et al. (2016), and is also known as the evolute. It governs the stability of the optimization problem over the real numbers. In contrast with the number of complex critical points, which is generically constant, the number of real critical points is constant only on the connected components of $\mathbb{R}^{n}\setminus\Sigma(\mathcal{V})$.
@@ 4.2 Parameter Discriminant
$\begin{cases}2r&\text{if }\text{rank}(M_{\bm{\theta}})=\text{rank}(A_{\bm{\theta}})+1\\ 2r+1&\text{if }\text{rank}(M_{\bm{\theta}})=\text{rank}(A_{\bm{\theta}})+2\\ 2r-2&\text{if }\text{rank}(M_{\bm{\theta}})=\text{rank}(A_{\bm{\theta}}).\end{cases}$
@@
The algebraic characterization of the parameter discriminant could be used for complexity-aware model selection. By Corollary 4.7, strata with strictly lower ED degree are defined by the vanishing of specific polynomials. Thus, we may formulate an algebraic regularizer that targets a specific complexity class by augmenting the loss with a penalty term. For instance, $\mathcal{L}(\bm{\theta})+\lambda(\Delta_{\text{circ}}(\bm{\theta}))^{2}$ would encourage models of lower algebraic complexity, trading generic expressivity for a decision boundary that is cheaper to verify.
@@ 5 Conclusion
We introduced an algebraic framework for robustness verification that connects neural network verification with tools from metric algebraic geometry. By characterizing verification through the Euclidean Distance degree and associated discriminants, our results provide an architecture-, data-, and parameter-dependent notion of intrinsic verification complexity, enabling the identification of decision boundaries and test instances that are provably easier or harder to verify.
Our approach complements existing verification methods by offering exact, structure-aware guarantees and by clarifying how architectural, parametric, and data-dependent properties govern verification difficulty. This perspective suggests new directions for algorithm and architecture design that reduce verification complexity.
Important open directions remain. Our analysis focuses on Euclidean distance, and alternative metrics may reveal different complexity phenomena. Moreover, while our framework applies directly to polynomial classifiers, extending algebraic complexity measures and exact certification techniques to broader piecewise-algebraic or non-polynomial networks is an important challenge for future work.
@@ Acknowledgments
A Numerical Solution via Homotopy Continuation
A.1 Formulation and the Gamma Trick
$H(\bm{z},t;\gamma):=\gamma(1-t)G(\bm{z})+tF(\bm{z}),$
A.2 Path Tracking via Predictor–Corrector Methods
A.3 Implementation and Post-Processing
B Proofs
B.1 Proof of Proposition˜2.1

### v2

@@ 3.1 ED Degree for Generic Parameters
Our framework allows us to distinguish between the geometric complexity of the boundary and the identifiability of the parameters. Let $p$ be the number of parameters and $N$ be the dimension of the space of degree-$D$ polynomials. Let $\Phi:\mathbb{R}^{p}\to\mathbb{R}^{N}$ be the parameterization map sending weights to coefficients. The observed drop in ED degree is governed by the dimension of the image $\mathrm{Im}(\Phi)$ relative to the ambient dimension $N$. For intermediate width architectures, we typically have $\dim(\mathrm{Im}(\Phi))\ll N$. This strict inequality ensures that the decision boundary lies on a proper subvariety of the polynomial space, which explains the reduced algebraic complexity. In contrast, parameter identifiability is determined by the dimension of the image relative to the domain dimension $p$. If $\dim(\mathrm{Im}(\Phi))<p$, then by the Fiber Dimension Theorem (Hartshorne, 1977, Chapter 1), the fibers $\Phi^{-1}(f)$ are positive-dimensional varieties, and thus the network is defective. This comparison clarifies the trade-off: the ED degree depends on how small the image is inside $\mathbb{R}^{N}$, while identifiability depends on how compressed the map is from $\mathbb{R}^{p}$. A network can thus achieve low verification complexity while remaining non-identifiable if the architecture is over-parameterized.
@@ 3.2 Expected Real ED Degree
The results in the preceding section concern critical points of the distance minimization problem over the complex field. This is the natural quantity for homotopy continuation, which computes all isolated complex solutions of the critical system. For robustness verification, however, one is often ultimately interested only in the real critical points, since these correspond to candidate nearest points on the decision boundary in the real input space. In practice, the number of real ED critical points is typically much smaller than the ED degree and may vary substantially across different generic realizations of the network parameters. To capture this phenomenon, we study the expected real ED degree under a distribution on the parameters.
A convenient tool for this purpose is the Kac–Rice formalism (Kac, 1943; Rice, 1944), which gives an exact expression for the expected number of zeros of a random field. Applied to our ED critical system, it yields a formula for the expected number of real critical points encountered in robustness verification.
@@
The above theorem is quite general since it applies to any sufficiently regular random field satisfying the stated nondegeneracy and integrability assumptions. The difficulty is not the applicability of Kac–Rice, but the evaluation of the resulting integral. Indeed, the formula involves the joint density of $(f(\bm{x}),\nabla f(\bm{x}))$ and a conditional expectation depending on the Hessian $\nabla^{2}f(\bm{x})$. For a generic random field, these quantities are typically hard to characterize explicitly.
To make the computation tractable, we therefore specialize to the neural network Gaussian process (NNGP) regime (Lee et al., 2018). This specialization is analytically convenient because, in the infinite-width limit, the random network induces a Gaussian field whose law is completely determined by its kernel. Consequently, the random variables $f(\bm{x})$, $\nabla f(\bm{x})$, and $\nabla^{2}f(\bm{x})$ are jointly Gaussian, so the density and conditional moments appearing in Theorem˜3.10 can be expressed in terms of kernel derivatives. Thus, while Kac–Rice itself is model-agnostic, the NNGP structure makes the formula sufficiently explicit to derive concrete bounds in high dimension and a closed-form expression in dimension one.
Theorem 3.11.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be the neural network Gaussian process (NNGP) induced by an infinite-width two-layer network with polynomial activation of degree $d\geq 1$. Then there exists a constant $C_{n,\bm{\xi}}>0$, independent of $d$, such that
$\mathbb{E}\Big[\#\{(\bm{x},\lambda)\in\mathbb{R}^{n}\times(\mathbb{R}\setminus\{0\}):F(\bm{x},\lambda)=\mathbf{0}\}\Big]\leq C_{n,\bm{\xi}}\,d^{n/2}.$
In particular, for fixed $n$, the expected real ED degree is $O(d^{n/2})$ as $d\to\infty$.
Proposition 3.12.
Let $f:\mathbb{R}\to\mathbb{R}$ be the neural network Gaussian process (NNGP) induced by an infinite-width two-layer network with polynomial activation of degree $d\geq 1$. Then the expected real ED degree is
$\mathbb{E}\Big[\#\{\,(x,\lambda)\in\mathbb{R}\times(\mathbb{R}\setminus\{0\}):F(x,\lambda)=\mathbf{0}\,\}\Big]=\frac{d}{\sqrt{2d-1}}.$
By Theorem˜3.11, for fixed $n$ the expected real ED degree is $O(d^{n/2})$, whereas the generic complex ED degree scales as $d^{n}$ by Proposition˜3.2. Hence the real critical-point count grows like the square root of the complex algebraic count. This is reminiscent of the classical Kostlan–Shub–Smale model for random homogeneous polynomial systems: for the orthogonally invariant Gaussian ensemble of degrees $d_{1},\dots,d_{n}$, the expected number of real projective roots is $\sqrt{d_{1}\cdots d_{n}}$, i.e., the square root of the generic complex Bézout count $d_{1}\cdots d_{n}$ (Kostlan, 1993; Edelman and Kostlan, 1995; Kostlan, 2002). Although the ED critical system arising from NNGPs is not a standard Kostlan system, it exhibits the same qualitative phenomenon: most complex critical points do not materialize over the reals, but the expected real count still grows polynomially with degree. In dimension one, Proposition˜3.12 gives the exact asymptotic $\frac{d}{\sqrt{2d-1}}\sim\sqrt{d/2}$.
@@ 4.1 ED Discriminant
The notion of ED discriminant was introduced in the context of algebraic optimization by Draisma et al. (2016), and is also known as the evolute. It governs the stability of the optimization problem over the real numbers. In contrast to the number of complex critical points, which is generically constant, the number of real critical points is constant only on the connected components of $\mathbb{R}^{n}\setminus\Sigma(\mathcal{V})$.
@@ 4.2 Parameter Discriminant
$\begin{cases}2r+1&\text{if }\text{rank}(M_{\bm{\theta}})=\text{rank}(A_{\bm{\theta}})+2\\ 2r&\text{if }\text{rank}(M_{\bm{\theta}})=\text{rank}(A_{\bm{\theta}})+1\\ 2r-2&\text{if }\text{rank}(M_{\bm{\theta}})=\text{rank}(A_{\bm{\theta}}).\end{cases}$
@@
The algebraic characterization of the parameter discriminant could be used for complexity-aware model selection. By Corollary 4.7, strata with strictly lower ED degree are defined by the vanishing of specific polynomials. Thus, we may formulate an algebraic regularizer that targets a specific complexity class by augmenting the loss with a penalty term. For instance, $\mathcal{L}(\bm{\theta})+\lambda(\Delta_{\text{sing}}(\bm{\theta}))^{2}$ would encourage models of lower algebraic complexity, trading generic expressivity for a decision boundary that is cheaper to verify. We illustrate this in Figure 5, where a quadratic network is trained on XOR data both with and without the algebraic regularizer encouraging $\Delta_{\text{sing}}(\bm{\theta})=0$. Even a modest penalty ($\lambda=0.1$) steers the decision boundary toward the degenerate locus ($x_{1}x_{2}=0$), reducing the ED degree from $4$ to $2$.
Figure 5: Left: decision boundary of a quadratic network trained on XOR data. Right: decision boundary of the same network trained with an algebraic regularizer that encourages $\Delta_{\text{sing}}(\bm{\theta})=0$.
5 Lightning Self-Attention
We now move beyond the feedforward polynomial network setting of the preceding sections and study these complexity measures numerically for lightning self-attention modules.
5.1 ED Degrees of Lightning Self-Attention
We study the ED degree of the decision boundary induced by lightning self-attention followed by a binary linear readout. Fix integers $t,d,d^{\prime},a$, and let
$\varphi_{W}(X)=VX(X^{\top}AX)\in\mathbb{R}^{d^{\prime}\times t},$
where $X\in\mathbb{R}^{d\times t}$ is the input with embedding dimension $d$ and sequence dimension $t$, and $W=(V,A)$ is the parameter with value matrix $V\in\mathbb{R}^{d^{\prime}\times d}$ and attention matrix $A\in\mathbb{R}^{d\times d}$ subject to a rank constraint $\operatorname{rank}(A)\leq a$. Thus the attention block produces a $d^{\prime}\times t$ output matrix whose entries are cubic polynomials in the $N=dt$ entries of $X$. Vectorizing this matrix gives a point in $\mathbb{R}^{d^{\prime}t}$, and composing with an affine map
$L:\mathbb{R}^{d^{\prime}t}\to\mathbb{R}^{2}$
produces two polynomial outputs
$f(X)=(f_{1}(X),f_{2}(X)).$
Since each coordinate of $\varphi_{W}(X)$ is cubic in the entries of $X$, each $f_{i}(X)$ is also a cubic polynomial in the $N=dt$ input variables. We define the decision boundary to be the hypersurface of points $X$ satisfying
$f_{1}(X)-f_{2}(X)=0.$
This is the natural boundary associated to the binary classifier, since the predicted label is determined by which of the two output scores is larger, and the prediction changes precisely when the two scores are equal. In particular, the polynomial $f_{1}-f_{2}$ records the locus of ambiguity between the two classes.
The ED degree of this hypersurface measures the number of complex critical points of the squared Euclidean distance from a generic point in the input space to the decision boundary. It therefore provides a concrete algebraic measure of the complexity of the nearest-boundary problem. Comparing this value to the ED degree of a generic cubic hypersurface in $\mathbb{R}^{N}$ reveals how far the lightning self-attention boundary is from generic cubic behavior. Table 2 summarizes the numerically computed ED degrees for these lightning self-attention decision boundaries, where $N=dt$ denotes the total number of input variables. Here $G$ denotes the ED degree of a generic cubic hypersurface in $\mathbb{R}^{N}$, which we use only as a coarse dense benchmark. Since lightning self-attention decision boundaries are highly structured and sparse, a more refined comparison should take their monomial support into account; see (Lindberg et al., 2023) for a general sparse optimization perspective in which critical-point counts are controlled by the associated sparse Lagrange system and mixed-volume data.
Table 2: ED degrees for lightning self-attention decision boundaries.
NN 4 6 8
$G$ 45 189 765
$t$ 4 1 2 2 2 2 6 1 3 2 3 3 2 2 2 2 3 2 2 8 1 4 2 4 2 2 2 2 2 2 2 2 2 2 2 2 2
$d$ 1 4 2 2 2 2 1 6 2 3 2 2 3 3 3 3 2 3 3 1 8 2 4 2 4 4 4 4 4 4 4 4 4 4 4 4 4
$d^{\prime}$ 1 1 1 2 1 2 1 1 1 1 1 2 2 3 1 1 2 2 2 1 1 1 1 2 2 3 4 1 1 1 2 3 4 2 3 4 3
$a$ 1 1 1 1 2 2 1 1 1 1 2 1 1 1 2 3 2 2 3 1 1 1 1 1 1 1 1 2 3 4 2 2 2 3 3 3 4
ED 7 12 17 28 33 39 7 12 17 20 39 46 49 49 67 83 99 132 143 7 12 17 20 50 56 56 56 85 141 157 359 359 359 468 468 468 479
NN 9 10
$G$ 1533 3069
$t$ 9 1 3 3 3 3 3 3 3 3 3 10 1 5 2 5 5 2 2 5 2 2 2 2 2 2 2
$d$ 1 9 3 3 3 3 3 3 3 3 3 1 10 2 5 2 2 5 5 2 5 5 5 5 5 5 5
$d^{\prime}$ 1 1 1 1 2 3 1 2 3 2 3 1 1 1 1 1 2 2 1 2 1 1 1 2 2 2 5
$a$ 1 1 1 2 1 1 3 2 2 3 3 1 1 1 1 2 1 1 2 2 3 4 5 2 3 4 5
ED 7 12 20 79 85 106 113 654 738 847 895 7 12 17 20 39 50 56 88 117 183 239 255 481 1034 1144 1155
Table 2 shows that the ED degree of the lightning self-attention decision boundary is substantially smaller than that of a generic cubic hypersurface in $\mathbb{R}^{N}$. Thus, although the boundary is cubic in $N=td$ variables, it is far from generic. In particular, the gap between the observed ED degrees and $G$ should be read not only as a sign of non-genericity, but also as evidence that the sparse and constrained algebraic structure of the attention boundary lowers the complexity of the nearest-boundary problem. The table suggests three basic trends:
  • Dependence on the factorization $N=td$. For fixed $N$, the ED degree depends strongly on the decomposition $N=td$ and generally grows faster with $d$ than with $t$. This reflects the asymmetric role of the embedding and sequence dimensions in the attention map $\varphi_{W}(X)=VX(X^{\top}AX)$.
  • Dimensional stabilization and cylindrical boundaries. For highly restricted architectures, the ED degree stabilizes as $N$ grows. For example, when $d^{\prime}=a=1$, it equals $12$ for $t=1$ and all $d\geq 2$, and $7$ for $d=1$ and all $t\geq 2$. This reflects the fact that low-rank weights force the boundary to depend on only a few linear combinations of the input, so the hypersurface behaves like a cylinder over a low-dimensional base.
  • Monotonicity and projection saturation. For fixed $t$ and $d$, the ED degree increases with the rank $a$ and the dimension $d^{\prime}$, but in our experiments it saturates once $d^{\prime}\geq t$. For example, when $t=2$ and $d=4$, increasing $d^{\prime}$ beyond $2$ does not change the ED degree. This is consistent with the constraint $\operatorname{rank}(\varphi_{W}(X))\leq t$, so projecting into dimensions beyond $t$ does not introduce additional geometric complexity.
Computation.
We compute the exact ED degrees using the HomotopyContinuation.jl package in Julia. A naive approach to this computation would fix the network weights randomly and track the full generic bound of $G$ paths to solve the critical equations. However, because our structured boundary has an ED degree strictly smaller than $G$, the majority of these generic paths diverge to infinity, causing numerical instability and artificially dropping valid solutions.
To bypass this, we use a specialized parameter homotopy via monodromy. Crucially, we treat the attention weights as parameters of the polynomial system alongside the target data point $u$:
@var X[1:d, 1:t]; x = vec(X)
@var K[1:d, 1:a]; @var Q[1:d, 1:a]; @var V[1:d_prime, 1:d]
@var W_out[1:2, 1:(d_prime*t)]; @var b_out[1:2]
@var u[1:(d*t)]; @var lambda
all_params = [vec(K); vec(Q); vec(V); vec(W_out); vec(b_out); u]
sys = System([x - u - lambda * grad_g; g], [x; lambda], all_params)
To initialize monodromy_solve, we construct an explicit seed solution. We sample a random complex input $x_{0}$, multiplier $\lambda_{0}$, and network weights, then choose the bias $b_{\text{out}}$ so that $g(x_{0})=0$ and set $u_{0}=x_{0}-\lambda_{0}\nabla g(x_{0})$. This yields a valid critical point for the corresponding parameter values:
# Calculate bias to satisfy g(x) = 0
b0_2 = randn(ComplexF64)
b0_1 = b0_2 - sum((W0[1, :] .- W0[2, :]) .* Y0_flat)
b0 = [b0_1, b0_2]
# Calculate u0 to trivially satisfy the critical equations
u0 = x0 - lambda0 * grad_g_eval
Starting from this seed, monodromy_solve explores the remaining solutions by following loops in parameter space. In practice, the default stopping heuristic stabilizes at the ED degree of the given architectural family without tracking the divergent paths of the generic cubic.
5.2 ED Discriminant for Lightning Self-Attention
Computing the ED discriminant symbolically by elimination is infeasible for the attention modules we consider, so we use numerical algebraic geometry instead. We focus on the architecture $(t,d,d^{\prime},a)=(2,2,1,1)$, for which the ambient data space is $\mathbb{R}^{2\times 2}$. We sample the network parameters from a standard normal distribution, thereby fixing a cubic hypersurface $g(X)=0$ with ED degree $17$, as established in Table 2. Hence a generic data point $u\in\mathbb{R}^{2\times 2}$ has exactly $17$ complex critical points for the distance function. Since non-real critical points occur in complex conjugate pairs, the number of real critical points is necessarily odd.
The ED discriminant is the real hypersurface in the data space where the number of real critical points changes. It partitions $\mathbb{R}^{2\times 2}$ into open chambers, with each chamber characterized by a constant, odd number of real critical points. To visualize these chambers, we restrict our view to random two-dimensional affine slices of the data space. We choose a generic base point $u_{0}\in\mathbb{R}^{2\times 2}$ and two orthogonal unit vectors $v_{1},v_{2}\in\mathbb{R}^{2\times 2}$, parameterizing a 2D plane as $u(s,c)=u_{0}+sv_{1}+cv_{2}$.
Number of real critical points:  $\blacksquare$ 3  $\blacksquare$ 5  $\blacksquare$ 7  $\blacksquare$ 9
Figure 6: Two-dimensional slices of the data space for a fixed self-attention architecture. The left and center panels display $6\times 6$ windows centered at the origin, while the right panel displays a $6\times 6$ window centered at (7,4)(-7,4) to highlight the 9-root chamber. Data points are color-coded by the number of real critical points.
To approximate these chambers, we sample 15,000 points $(s,c)$ in a bounding box. For each sample, we use parameter homotopy to track the 17 complex critical points from $u_{0}$ to the target $u(s,c)$, using HomotopyContinuation.jl. We treat the data point $u$ as the sole parameter of the critical system:
# grad_g is the gradient of the decision boundary g(x) = 0
# The variables are x and lambda; the data point u is the parameter
sys = System([x - u - lambda * grad_g; g], [x; lambda], u)
# Track the 17 base solutions to the random target parameter
res = solve(sys, base_sols; start_parameters = u0,
target_parameters = u_target)
Counting real solutions at the sampled points yields a discrete approximation to the chamber decomposition. Figure 6 shows three such slices, color-coded by the number of real critical points; the visible interfaces are numerical approximations to the intersections of the slices with the ED discriminant.
We include a brief discussion of the extension of this perspective to ReLU networks in Appendix H.
@@ 6 Conclusion
We introduced an algebraic framework for robustness verification that connects neural network verification with tools from metric algebraic geometry. By studying the Euclidean distance (ED) degree and the associated discriminants, we obtained intrinsic measures of verification complexity that depend on the architecture, the parameters, and the input data. This perspective makes it possible to analyze when verification is generically hard, when complexity drops, and how these changes are reflected in the geometry of the decision boundary.
This approach complements existing verification methods by providing exact, structure-aware guarantees and by clarifying how geometric and algebraic properties of a model govern verification difficulty. More broadly, it suggests that verification complexity can be studied as a mathematical invariant of a model class, rather than solely as a feature of a particular algorithm. The proposed framework may enable a systematic investigation of architectural features that reduce verification complexity, thereby informing the design of neural architectures and training procedures with favorable verification properties.
Several directions remain open. On the theoretical side, an important direction is to study the expected real ED degree for deep polynomial networks and attention architectures, as well as to derive analytic descriptions of the ED and parameter discriminants for specific polynomial models. In the case of lightning self-attention, we provided numerical evidence that the decision boundaries exhibit highly non-generic ED behavior. Establishing corresponding theoretical results (formulas, bounds, or structural explanations for the ED degrees and discriminants of attention modules) is a particularly promising direction for future work. On the computational side, a central challenge is to scale homotopy continuation methods to substantially larger architectures.
Regarding extensions, our analysis focuses on Euclidean distance, and it would be natural to study how the picture changes under alternative metrics or more general optimization objectives. Extending these algebraic complexity measures and exact certification techniques to broader non-polynomial or piecewise-algebraic architectures remains a major challenge.
@@ Acknowledgments
A Verification Problem
A.1 Proof of Proposition˜2.1